OpenStudy (anonymous):

The area of a circle is equal to the area of a square. Which one is greater? A. circumference of the circle B. perimeter of the square

OpenStudy (phi):

you can figure it out, but it is "well-known" (to those who know it well) that the shortest perimeter enclosing a fixed area is that of a circle.

OpenStudy (anonymous):

What does that mean?

OpenStudy (dumbcow):

it means the circle will always have the smallest perimeter compared to another shape

OpenStudy (anonymous):

so i don't have to go through steps to figure this out? The circle always has the smallest circumference/perimeter compared to every shape out there?

OpenStudy (phi):

if you did not know that, you start by saying \[ \pi r^2 = s^2 \\ s= r \sqrt{\pi} \] the perimeter of a circle vs square is \[ 2 \pi r \ vs \ 4s \\2 \pi r \ vs\ 4r \sqrt{\pi} \] divide both sides by 2r sqr(pi) \[ \sqrt{\pi} \ vs\ 2 \] pi is about 3.14 and its square root is less than the square root of 4 thus the perimeter of the circle is shorter

OpenStudy (dumbcow):

^^ what they said

OpenStudy (phi):

** compared to every shape out there?*** yes, that have the same area

OpenStudy (anonymous):

got it okay thanks phi!!!! <333